How many atoms are in a 591 g sample of gold


1
Expert's answer
2021-02-05T04:06:51-0500

mass of sample of gold=591gm

molar mass of gold=197g/mol

"mole= \tfrac{mass}{molar mass}\\

mole=\frac{591}{197}\\

mole= 3\\

Total number of atoms= mole* 6*10^{23}\\

Atoms=3*6*10^{23}\\

Total number of atoms= 18*10^{23}"